OpenStudy (anonymous):

This question to me seems to be missing info. If the force acting on an object stays the same, then the acceleration of the object is inversely proportional to its mass. If an object with a mass of 9 kilograms accelerates at a rate of per second per second by a force, find the rate of acceleration of an object with a mass of that is pulled by the same force.

OpenStudy (amistre64):

a = k/m

OpenStudy (amistre64):

since k = ma of the original situtation: k = 9*1 = 9 a = 9/m

OpenStudy (anonymous):

oh you just use 1

OpenStudy (amistre64):

if i can read it right, which i think i ran over a few details lol

OpenStudy (anonymous):

so then where do you go from there

OpenStudy (amistre64):

If the force acting on an object stays the same, then the acceleration of the object is inversely proportional to its mass: a=k/m If an object with a mass of 9 kilograms accelerates at a rate of per second per second by a force, at a rate of persec persec ??

OpenStudy (turingtest):

looks like it's missing info to me

OpenStudy (amistre64):

either a copy/paste went bad; or the original details is missing

OpenStudy (anonymous):

i think they messed up because that is how it is worded I thought something was missing to

OpenStudy (anonymous):

because you have to divide 9 by something right

OpenStudy (amistre64):

if you filled in the missing data with say "n" you could work out the rest :)

OpenStudy (amistre64):

n = k/9 such that the constant of variation = 9n

OpenStudy (amistre64):

find the rate of acceleration of an object with a mass of _??_ that is pulled by the same force.

OpenStudy (turingtest):

between the "rate of _____ per second per second" and "an object with a mass of _____ that is pulled" this thing is dead in the water

OpenStudy (anonymous):

but you would still have to do something with the 9 because my choices are a. 3 meters per second b. 27 c. 18 D. 24

OpenStudy (amistre64):

you cant determine a specific value without the missing data

OpenStudy (turingtest):

all you know is that the first object is\[a_1=\frac F9\]\[a_2=\frac F{m_2}\]which is not enough info to solve for\(a_2\) which is what they want you have three equations with two unknowns

OpenStudy (turingtest):

strike the "is the first object" part above

OpenStudy (turingtest):

sorry, I also meant two equations with three unknowns the typos never end

OpenStudy (anonymous):

yeah I can never get a hold of the stupid school so I guess I will just guess

OpenStudy (turingtest):

bummer... that happens too often in my opinion